OpenStudy (henrietepurina):

Can someone just help me show my work in algebra 1?

OpenStudy (alexandervonhumboldt2):

k

OpenStudy (henrietepurina):

Here is my question: (3x^3)y + 12xy - (3x^2)y - 12y Rewrite the expression completely factored. Show the steps of your work. And my answer was (3x^3)y−(3x^2)y+12xy−12y

OpenStudy (henrietepurina):

(refresh page to remove question marks)

OpenStudy (henrietepurina):

@asnaseer

OpenStudy (asnaseer):

Didn't we just do exactly this expression?

OpenStudy (henrietepurina):

yeah yeah, but in this one I need to factor it :P

OpenStudy (texaschic101):

3x^3y + 12xy - (3x^2y) - 12y --- factor out common terms...3 and y 3y(x^3 + 4x - x^2 - 4) = 3y (x^2 + 4)(x - 1)

OpenStudy (asnaseer):

Ok - what confused me is your statement: "And my answer was (3x^3)y−(3x^2)y+12xy−12y" which is not what we got last time

OpenStudy (asnaseer):

last time we got to:\[3y(x^3+4x-x^2-4)\]

OpenStudy (henrietepurina):

OMG.... I just messed up, my final answer is not (3x^3)y−(3x^2)y+12xy−12y :P

OpenStudy (asnaseer):

Ok, so look at where we got to:\[3y(x^3+4x-x^2-4)\]Now make use of your conclusion in your previous question which said that this will not change if we change the order of the terms inside the brackets. Therefore we can swap the positions of the \(4x\) and the \(-x^2\) to get:\[3y(x^3+4x-x^2-4)=3y(x^3-x^2+4x-4)\]Next we factor the first two terms in this to get:\[3y(x^3+4x-x^2-4)=3y(x^3-x^2+4x-4)=3y(x^2(x-1)+4x-4)\]Then factor the last two terms to get:\[=3y(x^2(x-1)+4(x-1))\]Now notice that both inner terms have a factor of \((x-1)\)

OpenStudy (asnaseer):

So we can factor out this \((x-1)\) to finally get:\[3y(x^3+4x-x^2-4)=3y(x-1)(x^2+4)\]

OpenStudy (henrietepurina):

Thanks again, you're a big help :)

OpenStudy (asnaseer):

yw :)